OpenStudy (anonymous):

f(x)=4x+7 find f(3) a. 15 c. 19 b. 12 d. 5

OpenStudy (callisto):

f(x)=4x+7 f(3)=4(3)+7 = ...

OpenStudy (anonymous):

thx